OpenStudy (anonymous):

Verify that f and g are inverse functions. f(x)=4x^2, x greater and equal to 0; g(x)=√x/2

OpenStudy (anonymous):

u begin by makeing f(x)=4x^2 into y=4x^2. then u swap x and y making it x=4y^2. then u begin solving for y

OpenStudy (anonymous):

\[y^2 = x/4 \] \[ y = \sqrt{x/4} = \sqrt{x}/2 \] therefore f(x) and g(x) are inverses

OpenStudy (anonymous):

i dont get that